# ZNorm: Z-Score Gradient Normalization for Deep Neural Networks

### v4

@@ Abstract
The rapid advancements in deep learning necessitate better training methods for deep neural networks (DNNs). As models grow in complexity, vanishing and exploding gradients impede performance. We propose Z-Score Normalization for Gradient Descent (ZNorm), an innovative technique that adjusts only the gradients to accelerate training and improve model performance. ZNorm normalizes the overall gradients, providing consistent gradient scaling across layers, thereby reducing the risks of vanishing and exploding gradients, having better performances. Our extensive experiments on CIFAR-10 and medical datasets demonstrate that ZNorm enhances performance metrics. ZNorm consistently outperforms existing methods, achieving superior results using the same experimental settings. In medical imaging applications, ZNorm improves tumor prediction and segmentation performances, underscoring its practical utility. These findings highlight ZNorm’s potential as a robust and versatile tool for enhancing the training speed and effectiveness of deep neural networks across a wide range of architectures and applications.
@@ 1 Introduction
The rapid advancements in deep learning have revolutionized various fields, from computer vision to natural language processing, driven by the availability of large-scale datasets, increased computational power, and the development of sophisticated network architectures. However, training deep neural networks (DNNs) [9] remains a formidable challenge. As models become deeper and more complex, issues such as vanishing and exploding gradients frequently arise, impeding model performance [11].
To address these challenges, several normalization techniques have been proposed to improve the training process. Methods such as batch normalization (BN) [5], layer normalization (LN) [1], and weight standardization (WS) [15] have been widely adopted for their effectiveness in stabilizing training by adjusting the mean and variance of activations or weights. Another promising recent approach, Gradient Centralization [22], focuses on normalizing the gradients themselves by adjusting their mean to zero, which has shown to improve training consistency and performance. Similarly, Gradient Clipping [12] has been employed to prevent the exploding gradient problem by capping gradients at a predefined threshold. Weight decay [8], another widely used technique, applies L2 regularization to the weights during training to prevent overfitting and control weight growth, which indirectly improves the stability of the training process.
Building on these approaches, our work aims to demonstrate that by simply normalizing and adjusting gradients, significant improvements in performance can be achieved across a variety of network architectures. We introduce ZNorm, a technique designed to normalize gradients, providing consistent gradient scaling. This enhances the generalization capabilities of DNNs. Unlike other methods that primarily focus on activation or weight normalization, ZNorm directly addresses the gradients, leading to faster training. Instead of striving to develop increasingly complex architectures, our approach suggests that substantial performance gains can be realized by refining the training process itself. By improving how gradients are handled during training, we achieve better results with existing architectures, thereby challenging the notion that more complex models are always necessary for better performance.
Figure 1: Visualization of the ZNorm($\Phi$) process applied to gradient matrices and tensors in both fully-connected and convolutional layers. The process includes calculating the gradient mean vector and standard deviation vector, followed by normalizing the gradients via ZNorm.
The impact of ZNorm is particularly noteworthy as it achieves these performance gains, making it an accelerator for improving training processes. Our extensive experiments on both synthetic and real-world datasets validate the effectiveness of ZNorm. The results consistently show that ZNorm boosts performance metrics, leading to more accurate models. For instance, in medical imaging applications, ZNorm has demonstrated its practical utility by improving the accuracy of tumor prediction and segmentation tasks. This makes ZNorm a valuable tool for advancing deep learning models across a wide range of architectures and applications by focusing on the crucial aspect of gradient normalization.
@@ 2 Related Works
Optimization techniques are central to training deep neural networks effectively. Stochastic gradient descent (SGD) and its variants, such as SGD with momentum [14] and adaptive methods like Adam [6], have been crucial in optimizing complex architectures. These methods aim to minimize the loss function via parameter updates. Adaptive optimizers like Adagrad [2] and RMSProp [19] further adjust learning rates dynamically based on gradient magnitude and history, stabilizing the training process for deep models.
Normalization techniques help mitigate training instability and smooth out optimization landscapes. Batch normalization (BN) [5] normalizes activations across mini-batches to reduce internal covariate shifts. Layer normalization (LN) [1] and group normalization (GN) [21] normalize along different dimensions, such as layers or groups of channels, while weight normalization (WN) [17] and weight standardization (WS) [15] ensure consistent weight scaling, contributing to faster and more stable training.
Gradient adjustment methods directly manipulate the gradient updates to improve training stability and prevent issues like exploding or vanishing gradients. Gradient Clipping [13] is used to prevent the exploding gradient problem by capping the gradient magnitudes. If the gradient exceeds a predefined threshold $\tau$, it is rescaled like $\nabla_{\theta}L(\theta_{t})\leftarrow\nabla_{\theta}L(\theta_{t})\cdot\min\left(1,\frac{\tau}{||\nabla_{\theta}L(\theta_{t})||}\right)$. This ensures that no gradient update is too large, helping stabilize training in recurrent or deep networks.
@@
AdamW [10] improves upon traditional weight decay by decoupling L2 regularization from the gradient update in adaptive optimization methods like Adam. In traditional Adam, weight decay is applied implicitly through the L2 regularization term, which affects both the gradient and the parameter updates. AdamW, however, applies weight decay directly to the weights during the parameter update step, ensuring that the regularization only affects the weight magnitudes and not the gradient direction. The gradient update for weight decay in AdamW is modified as $\nabla_{\theta}L_{\text{decay}}(\theta_{t})=\nabla_{\theta}L(\theta_{t})+\lambda\theta_{t}$ where $\lambda$ is the weight decay coefficient. This decoupled approach in AdamW prevents the regularization term from interfering with the adaptive learning rate updates, leading to more stable training and better generalization, especially in large-scale models. Gradient Centralization [22] normalizes the gradients by subtracting their mean, ensuring a zero-mean gradient. The gradient update is adjusted as $\nabla_{\theta}L(\theta_{t})\leftarrow\nabla_{\theta}L(\theta_{t})-\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta_{i}}L(\theta_{t})$ where $n$ is the number of gradient elements. This technique improves generalization and speeds up convergence by regularizing the training dynamics.
Our proposed Z-Score Normalization (ZNorm) builds on these techniques. By applying Z-score normalization directly to overall gradients, we standardize gradient scaling across layers, reducing vanishing and exploding gradients [12] while ensuring consistent learning progress. This leads to improved generalization and robustness of the models.
@@ 3 Methodology
3.1 Notations
Consider a deep neural network consisting of $L$ layers, where each layer $l$ has associated weights $\mathbf{\theta}^{(l)}$. Fully connected layers can be expressed as $\mathbf{\theta}^{(l)}_{fc}\in\mathbb{R}^{D_{l}\times M_{l}}$, where $D_{l}$ represents the number of neurons and $M_{l}$ represents the dimension of the input to the layer. Convolutional layers, $\mathbf{\theta}^{(l)}_{conv}$, can be represented as a 4-dimensional tensor $\mathbf{\theta}^{(l)}_{conv}\in\mathbb{R}^{C_{out}^{(l)}\times C_{in}^{(l)}\times k_{1}^{(l)}\times k_{2}^{(l)}}$, where $C_{in}^{(l)}$ and $C_{out}^{(l)}$ are the number of input and output channels, respectively, and $k_{1}^{(l)}$ and $k_{2}^{(l)}$ represent the kernel sizes. Let $\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ denote the gradient of the loss function $\mathcal{L}$ with respect to the weights $\mathbf{\theta}^{(l)}$ of layer $l$. $\nabla\mathcal{L}(\mathbf{\theta})$ denotes the overall gradient tensor.
3.2 Z-Score Normalization for Gradients
Z-Score Normalization is applied layer-wise, meaning each layer’s gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is individually normalized using ZNorm, which can be expressed as:
@@ 3.3 Embedding ZNorm to Adam
Z-Score Normalization (ZNorm) can be seamlessly integrated into existing deep neural network (DNN) optimization algorithm, Adam [6]. ZNorm, applied directly to the gradient tensors. In this paper, we specifically focus on embedding ZNorm into the Adam optimization algorithm, which is one of the most widely used optimization algorithms in deep learning due to its adaptive learning rate properties and effectiveness in training deep models. Algorithm 1 shows how ZNorm can be embedded into the Adam algorithm with minimal changes. The ZNorm step is applied immediately after computing the gradient, and then the normalized gradient is utilized in the subsequent Adam updates.
4 Theoretical Analysis
Z-score normalization normalizes gradients by subtracting their mean and dividing by the standard deviation, potentially altering the direction of individual components. This raises concerns about whether the overall descent direction is maintained. This analysis explains why, despite these changes, the optimization process remains effective. Specifically, it shows that the inner product of the normalized gradient with the original gradient remains positive, preserving the overall descent direction necessary for reducing the objective function.
Lemma 4.1. Let $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\in\mathbb{R}^{d}$ be the original gradient vector with mean $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}=\frac{1}{d}\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{i}^{(l)}})$, variance $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}={\frac{1}{d}\sum_{i=1}^{d}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})})^{2}}$. The Z-score normalized gradient $\Phi_{Z}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))$ is defined as
@@ 4 Theoretical Analysis
$\Phi_{Z}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))=\frac{\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}$
for $\text{for }i=1,2,\dots,d.$
Then, the inner product of the normalized gradient $\Phi_{Z}$ with the original gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is given by
@@
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla}\mathcal{L}(\mathbf{\theta^{(l)}})^{2}}{\sigma_{\nabla}\mathcal{L}(\mathbf{\theta^{(l)}})},$
where $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}$ is the Euclidean norm of the vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$. This result indicates that while the Z-score standardization can alter the direction of individual components of the gradient, the overall descent direction is preserved.
Proof. We begin by recalling the definition of the Z-score standardized gradient:
$\Phi_{Z}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))=\frac{\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
where $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}=\frac{1}{d}\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\quad\text{and}\quad\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}=\frac{1}{d}\sum_{i=1}^{d}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})})^{2}$ Here, $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ represents the mean of the gradient components, and $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ represents the standard deviation, which measures the spread of the gradient components around the mean.
Next, the z-score normalized gradient vector $\Phi_{Z}$ can be expressed in vector form as:
$\Phi_{Z}=\frac{1}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}(\nabla\mathcal{L}(\mathbf{\theta^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}\mathbf{1}),$
where $\mathbf{1}$ is the $d$-dimensional vector of ones.
Here, the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, which is given by:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\left(\frac{(\nabla\mathcal{L}(\mathbf{\theta^{(l)}})-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}\mathbf{1})}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}\right)^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}}).$
Expanding this expression, we get:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ $=\frac{((\nabla\mathcal{L}(\mathbf{\theta^{(l)}}))^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}$ (2)
$-\frac{-\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}\mathbf{1}^{T}{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})})}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}$ (3)
Now, we calculate $\mathbf{1}^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, which represents the sum of all components of the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$:
$\mathbf{1}^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})=d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}.$
Substituting this into the expression for $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, we obtain:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{(\nabla\mathcal{L}(\mathbf{\theta^{(l)}}))^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
Next, let’s express $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ in terms of the Euclidean norm of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\sum_{i=1}^{d}(\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}}))^{2}=\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}.$
Thus, the inner product becomes:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
This equation shows that the inner product of the Z-score normalized gradient $\Phi_{Z}$ with the original gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is determined by the difference between the squared Euclidean norm of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ and the scaled mean squared value of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, divided by the standard deviation $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$.
This result is significant because it indicates that while the Z-score standardization can change the direction of individual components of the gradient, the overall directionality, as measured by the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, remains aligned with the original gradient direction. Specifically, if $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is a descent direction, then $(\Phi_{Z})^{T}$ will also be a descent direction, as the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ remains positive, which proves the lemma.
Theorem 4.2. Given the same conditions as in the Lemma, the Z-score normalized gradient $\hat{g}$ preserves the descent direction of the original gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$. Specifically, if $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is a descent direction, then $(\Phi_{Z})$ is also a descent direction, i.e.,
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})>0.$
This implies that even though Z-score standardization may change the directions of individual gradient components, the overall gradient vector still points in a direction that will reduce the objective function.
Proof. From the result of the Lemma, we have:
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}.$
Here, $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}=\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})^{2}$ is the squared Euclidean norm of the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, and $\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ is the mean of the gradient components.
Since $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ is the standard deviation of the elements of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$, we have $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}>0$. The standard deviation $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}$ is positive as long as the components of $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ are not all identical (i.e., $g_{i}$ are not constant for all $i$).
Next, consider the term $d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}$. This term can be rewritten as:

### v5

# ZNorm: Z-Score Gradient Normalization Accelerating Skip-Connected Network Training without Architectural Modification

@@ Abstract
The rapid advancements in deep learning necessitate better training methods for deep neural networks (DNNs). As models grow in complexity, vanishing and exploding gradients impede performance, particularly in skip-connected architectures like Deep Residual Networks. We propose Z-Score Normalization for Gradient Descent (ZNorm), an innovative technique that adjusts only the gradients without modifying the network architecture to accelerate training and improve model performance. ZNorm normalizes the overall gradients, providing consistent gradient scaling across layers, effectively reducing the risks of vanishing and exploding gradients and achieving superior performance. Extensive experiments on CIFAR-10 and medical datasets confirm that ZNorm consistently outperforms existing methods under the same experimental settings. In medical imaging applications, ZNorm significantly enhances tumor prediction and segmentation accuracy, underscoring its practical utility. These findings highlight ZNorm’s potential as a robust and versatile tool for enhancing the training and effectiveness of deep neural networks, especially in skip-connected architectures, across various applications.
@@ 1 Introduction
Efficient training of deep neural networks (DNNs) has become increasingly critical as deep learning models continue to grow in depth and complexity[15]. However, training DNNs remains a challenging task due to issues such as vanishing and exploding gradients, which limit performance and scalability[18]. Recently, Vision Transformers (ViTs) [5] have gained popularity for their impressive performance in computer vision tasks. However, convolutional neural networks (CNNs), especially skip-connected networks like ResNet, DenseNet, and U-Net, remain widely used due to their effectiveness and robustness, particularly in scenarios with smaller datasets [16] or specific tasks like medical imaging [14, 31, 29]. Ensuring efficient training of skip-connected networks is crucial, as these architectures continue to play a significant role in a variety of applications.
Numerous methods have been developed to address gradient-related challenges [26], focusing primarily on normalizing activations or weights to stabilize training. Techniques such as Batch Normalization (BN)[10], Layer Normalization (LN)[1], and Weight Standardization (WS)[22] have become standard in many architectures. While these methods improve training stability, they often introduce architectural modifications, increasing implementation complexity and reducing flexibility in certain scenarios. Gradient-based techniques such as Gradient Centralization[32], Gradient Clipping[19], and Weight Decay[13] demonstrate that focusing on gradients alone can yield significant performance improvements without altering network structures.
Building upon this principle, we propose Z-score Gradient Normalization (ZNorm), a novel approach to efficient training that directly adjusts gradients to ensure consistent scaling across layers. By normalizing both the mean and variance of gradients, where stable gradient flow is critical to achieving optimal performance. Our core contributions in this paper are as follows:
  • We improve model performance by adjusting gradients only, without modifying neural network architecture.
  • ZNorm is simple to implement, making it adaptable to various models with skip-connection.
ZNorm exemplifies efficient training by accelerating training and improving model performance without requiring architectural changes. Unlike activation-based normalization techniques, ZNorm operates exclusively on gradients, making it a lightweight yet powerful tool for optimizing skip-connected networks like ResNet and U-Net. Its compatibility with these architectures enhances their robustness against vanishing and exploding gradients, enabling consistent and efficient training.
To demonstrate its effectiveness, we evaluate ZNorm on a variety of skip-connected networks across synthetic and real-world datasets. Experiments show that ZNorm consistently outperforms existing methods, achieving superior performance in tasks such as image classification and medical imaging applications. In particular, ZNorm significantly enhances tumor prediction and segmentation accuracy, showcasing its utility in high-impact scenarios. These findings establish ZNorm as a robust and versatile tool for efficient training of skip-connected networks, capable of improving both training speed and model performance across diverse applications.
@@ 2 Related Works
Optimization techniques are central to training deep neural networks effectively. Stochastic gradient descent (SGD) and its variants, such as SGD with momentum [21] and adaptive methods like Adam [11], have been crucial in optimizing complex architectures. These methods aim to minimize the loss function via parameter updates. Adaptive optimizers like Adagrad [6] and RMSProp [28] further adjust learning rates dynamically based on gradient magnitude and history, stabilizing the training process for deep models. AdamW [17] improves upon traditional weight decay by decoupling L2 regularization from the gradient update in adaptive optimization methods like Adam. In traditional Adam, weight decay is applied implicitly through the L2 regularization term, which affects both the gradient and the parameter updates. AdamW, however, applies weight decay directly to the weights during the parameter update step, ensuring that the regularization only affects the weight magnitudes and not the gradient direction. The gradient update for weight decay in AdamW is modified as $\nabla_{\theta}L_{\text{decay}}(\theta_{t})=\nabla_{\theta}L(\theta_{t})+\lambda\theta_{t}$ where $\lambda$ is the weight decay coefficient. This decoupled approach in AdamW prevents the regularization term from interfering with the adaptive learning rate updates, leading to more stable training and better generalization, especially in large-scale models.
Training deep neural networks (DNNs) effectively requires overcoming challenges such as vanishing and exploding gradients, which hinder the performance of complex models [18]. Traditional approaches to address these issues include normalization techniques applied to activations and weights, such as Batch Normalization (BN) [10], Layer Normalization (LN) [1], and Weight Standardization (WS) [22]. While effective, these techniques require modifications to the network architecture, which can add complexity and limit flexibility. Additionally, large-batch optimization techniques, such as LARS [34] and LAMB [33], provide layer-wise adaptive learning rates, enabling stable training with large batches. While these methods effectively scale learning rates based on weight norms, they are primarily designed for large-batch settings and may not generalize to smaller batch sizes or different architectures. In contrast, ZNorm enhances training stability through direct gradient normalization, eliminating the need for complex adjustments or specific batch configurations. Unlike batch-dependent methods, ZNorm performs normalization on the gradient tensors themselves, ensuring consistent behavior regardless of the chosen batch size.
Gradient adjustment methods directly manipulate the gradient updates to improve training stability and prevent issues like exploding or vanishing gradients. Gradient Clipping [19] is one such method that caps gradient magnitudes to a threshold $\tau$, effectively preventing the exploding gradient problem. If the gradient exceeds $\tau$, it is rescaled like $\nabla_{\theta}L(\theta_{t})\leftarrow\nabla_{\theta}L(\theta_{t})\cdot\min\left(1,\frac{\tau}{||\nabla_{\theta}L(\theta_{t})||}\right)$. While Gradient Clipping is effective for stabilizing training, it lacks comprehensive control over gradient distribution across layers, potentially leading to inconsistencies in gradient scaling. Gradient Centralization (GC) [32] is another method that normalizes gradients by zero-centering them, effectively enhancing training consistency and performance. GC achieves this by adjusting the gradient mean to zero $\nabla_{\theta}L(\theta_{t})\leftarrow\nabla_{\theta}L(\theta_{t})-\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta_{i}}L(\theta_{t})$ where $n$ is the number of gradient elements. Although GC improves generalization and speeds up convergence, it only adjusts the mean, leaving the variance uncontrolled, which may not fully stabilize training in deep networks. SING [4] is another recent approach, offering plug-and-play gradient standardization to improve stability with optimizers such as AdamW. SING normalizes gradients layer-wise, enhancing generalization and allowing escape from narrow local minima. However, SING’s layer-wise adjustments are dependent on specific optimizers and introduce additional hyperparameters, whereas ZNorm operates independently of optimizer choice and requires no extra tuning, making it simpler to implement.
Building on these approaches, our work introduces Z-score Gradient Normalization (ZNorm), a technique that standardizes both the mean and variance of gradients across layers, ensuring consistent scaling. By normalizing the gradients using Z-scores, ZNorm directly addresses both vanishing and exploding gradients, particularly enhancing stability in skip-connected architectures like ResNet and U-Net. Unlike GC, which only zero-centers the gradient, ZNorm performs full normalization $\nabla_{\theta}L(\theta_{t})\leftarrow\frac{\nabla_{\theta}L(\theta_{t})-\mu_{\nabla}}{\sigma_{\nabla}}$ where $\mu_{\nabla}$ and $\sigma_{\nabla}$ are the mean and standard deviation of the gradient elements, respectively. This additional variance control provides more stable training across diverse skip-connected architectures and applications.
In summary, ZNorm positions itself as a straightforward and adaptable gradient normalization technique that avoids architectural modifications. By focusing solely on gradient adjustments, ZNorm can be easily integrated into various training processes with minimal changes. This layer-wise gradient scaling achieves stable and efficient training, particularly enhancing performance in skip-connected networks such as ResNet, DenseNet, and U-Net. As a result, ZNorm proves to be a valuable tool for improving training speed and effectiveness across diverse applications.
Furthermore, by only adjusting the gradients, ZNorm simplifies the training pipeline, making it suitable for rapid implementation without the need for extensive reconfiguration or additional hyperparameters. This adaptability allows ZNorm to be applied flexibly across different network architectures and tasks, ensuring that stable training benefits are accessible without requiring fundamental changes to the underlying model design.
@@ 3 Methodology
Figure 1: Visualization of the ZNorm($\Phi$) process applied to gradient matrices and tensors in both fully-connected and convolutional layers. The process includes calculating the gradient mean vector and standard deviation vector, followed by normalizing the gradients via ZNorm. The training process showing how normalized gradients $\Phi(\nabla L(\theta))$ are integrated into the forward-backward optimization loop, with gradient adjustment occurring between gradient computation and the optimizer update step.
Preliminaries. Consider a deep neural network consisting of $L$ layers, where each layer $l$ has associated weights $\mathbf{\theta}^{(l)}$. Fully connected layers can be expressed as $\mathbf{\theta}^{(l)}_{fc}\in\mathbb{R}^{D_{l}\times M_{l}}$, where $D_{l}$ represents the number of neurons and $M_{l}$ represents the dimension of the input to the layer. Convolutional layers, $\mathbf{\theta}^{(l)}_{conv}$, can be represented as a 4-dimensional tensor $\mathbf{\theta}^{(l)}_{conv}\in\mathbb{R}^{C_{out}^{(l)}\times C_{in}^{(l)}\times k_{1}^{(l)}\times k_{2}^{(l)}}$, where $C_{in}^{(l)}$ and $C_{out}^{(l)}$ are the number of input and output channels, respectively, and $k_{1}^{(l)}$ and $k_{2}^{(l)}$ represent the kernel sizes. Let $\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ denote the gradient of the loss function $\mathcal{L}$ with respect to the weights $\mathbf{\theta}^{(l)}$ of layer $l$. $\nabla\mathcal{L}(\mathbf{\theta})$ denotes the overall gradient tensor.
Z-Score Normalization for Gradients. Z-Score Normalization is applied layer-wise, meaning each layer’s gradient $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is individually normalized using ZNorm, which can be expressed as:
@@ 3.1 Embedding ZNorm to Adam
Z-Score Normalization (ZNorm) can be seamlessly integrated into existing deep neural network (DNN) optimization algorithm, Adam [11]. ZNorm, applied directly to the gradient tensors. In this paper, we specifically focus on embedding ZNorm into the Adam optimization algorithm, which is one of the most widely used optimization algorithms in deep learning due to its adaptive learning rate properties and effectiveness in training deep models.
Algorithm 1 shows how ZNorm can be embedded into the Adam algorithm with minimal changes. The ZNorm step is applied immediately after computing the gradient, and then the normalized gradient is utilized in the subsequent Adam updates.
Stability of ZNorm Training. Gradient Centralization [32] has noted potential instability when applying Z-score standardization directly to gradients. Through our analysis, we identify that this instability primarily stems from the diminishing gradient phenomenon during training convergence.
First, as training progresses towards convergence, the loss function $\mathcal{L}(\mathbf{\theta})$ monotonically decreases, leading to progressively smaller gradients. This relationship can be formally expressed as:
$\|\nabla\mathcal{L}(\mathbf{\theta}^{(t+1)})\|\leq\|\nabla\mathcal{L}(\mathbf{\theta}^{(t)})\|$ (2)
where t denotes the training step [9, 24, 2].
As the model converges to a local minimum $\mathbf{\theta}^{*}$, the gradient approaches zero:
$\lim_{t\to\infty}\nabla\mathcal{L}(\mathbf{\theta}^{(t)})=\nabla\mathcal{L}(\mathbf{\theta}^{*})=0$ (3)
This convergence behavior creates a critical challenge for Z-score normalization. As gradients become extremely small, their standard deviation $\sigma_{\nabla\mathbf{\theta}}$ also approaches zero. Consequently, when normalizing the gradients:
$\frac{\nabla\mathcal{L}(\mathbf{\theta})-\mu_{\nabla\mathbf{\theta}}}{\sigma_{\nabla\mathbf{\theta}}}$ (4)
the denominator becomes vanishingly small, potentially causing the normalized gradients to explode. Even with a small epsilon term added for numerical stability, this fundamental issue persists near convergence.
Role of Skip Connection for ZNorm. The instability problem mentioned earlier can be resolved by using skip connections [7]. Skip connections help prevent gradients from becoming extremely small, thereby maintaining the stability of ZNorm during training. Skip connections are commonly used in various network architectures such as ResNet [7], DenseNet [8], and U-Net [23, 36]. Let’s consider a deep neural network without skip connections. For a layer $l$, the gradient of the loss with respect to the weights $\mathbf{\theta}^{(l)}$ can be expressed as:
@@
$\nabla\mathcal{L}(\mathbf{\theta}^{(l)})=\frac{\partial\mathcal{L}}{\partial\mathbf{\theta}^{(l)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{z}^{(l)}}\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}=\delta^{(l)}\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$ (5)
where $\delta^{(l)}$ is the error term at layer $l$, and $\mathbf{z}^{(l)}$ is the pre-activation (weighted input) at layer $l$ [7, 15]. For a deep network, this is recursively expanded using the chain rule:
@@
$\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ $=\delta^{(L)}\cdot\frac{\partial\mathbf{z}^{(L)}}{\partial\mathbf{z}^{(L-1)}}\cdots\frac{\partial\mathbf{z}^{(l+1)}}{\partial\mathbf{z}^{(l)}}\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$ (6)
$=\delta^{(L)}\cdot\prod_{k=l+1}^{L}\frac{\partial\mathbf{z}^{(k)}}{\partial\mathbf{z}^{(k-1)}}\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$ (7)
Each term $\frac{\partial\mathbf{z}^{(k+1)}}{\partial\mathbf{z}^{(k)}}$ represents the derivative of the pre-activation of layer $k+1$ with respect to the pre-activation of the previous layer $k$, which can be very small for deep networks, leading to gradient vanishing.
Consider the inclusion of skip connections in the form:
$\mathbf{z}^{(l)}=f(\mathbf{z}^{(l-1)})+\mathbf{z}^{(l-1)}$ (8)
where $f(\mathbf{z}^{(l-1)})$ is the activation function [7]. With skip connections, the gradient with respect to $\mathbf{\theta}^{(l)}$ at layer $l$ can be expressed as:
$\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{z}^{(l)}}\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$ (9)